# OLATverse: A Large-scale Real-world Object Dataset with Precise Lighting Control

### v1

@@ 4 Application
Figure 8: Visual comparison of pseudo ground truth normals with normals estimated by DR [30], RGBX [61], SN [58] and GW [15]. To ensure a robust and generalized comparison, we provide input images of each validation object under four different illuminations.
@@ 4.2 Inverse Rendering and View Synthesis
To construct a comprehensive validation dataset, we carefully select $42$ objects with rich textures spanning $14$ material categories. For each object, the raw images are downsampled to $750\times 1.4\text{k}$ resolution, and every fifth camera and light are used as inference sets, while the remaining data are utilized for training. Object meshes are reconstructed using the feature-based algorithm provided in MetaShape [1]. We evaluate Mitsuba [39] (using MetaShape-reconstructed meshes) as well as several 3DGS-based inverse rendering methods [3, 14, 34] on our dataset. All baseline methods are executed using the source code released by the authors and initialized with the same reconstructed meshes. We visualize relighting results under inference views and lighting conditions in  Fig. 7, and report quantitative metrics (SSIM [56], PSNR, and LPIPS [65]) in  Tab. 2. As shown, GS3 consistently outperforms other inverse rendering methods both visually and numerically. Compared to the other methods, GS3 accurately captures the specular reflections on glossy surfaces such as the watermelon, metallic rabbit, and plastic roadblock.
@@ 4.3 Normal Estimation
Next, we benchmark several diffusion-based normal estimation methods [30, 61, 58, 15] on our validation dataset. In OLATverse, the extracted surface normals from $5$ polarized views and $35$ non-polarized views are in the world coordinate. During preprocessing, we select normal maps extracted from $5$ polarized views for evaluation and reproject the surface normals from world to camera coordinates using camera parameters, ensuring compatibility with the monocular normal estimation methods. To enable lighting-invariant evaluation, each object is evaluated under four different environmental illuminations. For all experiments, we use official source code, and present both qualitative and quantitative results in  Fig. 8 and  Tab. 3. We follow the same evaluation protocol as StableNormal (SN) [58], reporting the mean and median angular errors, where the lower values indicates better accuracy. We also measure the percentage of pixels with angular error below thresholds of $11.25^{\circ}$, $22.5^{\circ}$, and $30^{\circ}$. Quantitatively, SN and GeoWizard (GW), which are specifically designed for normal estimation tasks, outperform RGBX and DiffusionRender(DR), which are originally developed for image/video relighting. Visually RGBX and DR produce the surface normals with finer high-frequency details than SN and GW. However, none of these methods recover accurate surface normals for real-world objects with complex geometric structures in the evaluation dataset, highlighting the importance of OLATverse for advancing normal estimation research.
@@
Mitsuba+Mshape [39] 35.906 0.0260 0.976
GS3 [3] 38.538 0.0263 0.982
RNG [14] 32.065 0.051 0.962
BiGS [34] 32.98 0.0451 0.940
@@
SN [58] 31.85 30.25 8.93 34.00 55.40
RGBX [61] 51.95 49.70 6.40 22.80 35.85
DR [30] 34.88 33.28 8.13 31.00 50.15
GW [15] 34.42 32.03 10.98 34.10 50.05
@@ 5 Limitation
Although we apply a linear polarized filter to cancel out the specular reflection during normal extraction, artifacts remain noticeable for objects with glossy materials or low-reflectance texture, consistent with the observation in RNHA [68]. These artifacts arise from the weak signal-to-noise ratio of low-reflectance objects and the ambiguity between view-dependent reflection and surface normals. While the extracted surface normals and diffuse albedo are not exact ground truth, they can still provide valuable supervision signals for multi-modal training tasks. In addition, the ground truth meshes are not incorporated in our dataset due to hardware limitations. In the future, it would be interesting to integrate an advanced scanning system to jointly capture both appearance and geometry of real objects.

### v2

@@ 4 Application
Figure 8: Visual comparison of pseudo ground truth normals with normals estimated by DR [liang2025diffusion], RGBX [zeng2024rgb], SN [ye2024stablenormal] and GW [fu2024geowizard]. To ensure a robust and generalized comparison, we provide input images of each validation object under four different illuminations.
@@ 4.2 Inverse Rendering and View Synthesis
To construct a comprehensive validation dataset, we carefully select $42$ objects with rich textures spanning $14$ material categories. For each object, the raw images are downsampled to $750\times 1.4\text{k}$ resolution, and every fifth camera and light are used as inference sets, while the remaining data are utilized for training. Object meshes are reconstructed using the feature-based algorithm provided in MetaShape [agisoft_metashape]. We evaluate Mitsuba [nimier2019mitsuba] (using MetaShape-reconstructed meshes) as well as several 3DGS-based inverse rendering methods [bi2024gs3, fan2025rng, liu2024bigs] on our dataset. All baseline methods are executed using the source code released by the authors and initialized with the same reconstructed meshes. We visualize relighting results under inference views and lighting conditions in  Fig. 7, and report quantitative metrics (SSIM [wang2004image], PSNR, and LPIPS [zhang2018unreasonable]) in  Tab. 2. As shown, GS3 consistently outperforms other inverse rendering methods both visually and numerically. Compared to the other methods, GS3 accurately captures the specular reflections on glossy surfaces such as the watermelon, metallic rabbit, and plastic roadblock.
@@ 4.3 Normal Estimation
Next, we benchmark several diffusion-based normal estimation methods [liang2025diffusion, zeng2024rgb, ye2024stablenormal, fu2024geowizard] on our validation dataset. In OLATverse, the extracted surface normals from $5$ polarized views and $35$ non-polarized views are in the world coordinate. During preprocessing, we select normal maps extracted from $5$ polarized views for evaluation and reproject the surface normals from world to camera coordinates using camera parameters, ensuring compatibility with the monocular normal estimation methods. To enable lighting-invariant evaluation, each object is evaluated under four different environmental illuminations. For all experiments, we use official source code, and present both qualitative and quantitative results in  Fig. 8 and  Tab. 3. We follow the same evaluation protocol as StableNormal (SN) [ye2024stablenormal], reporting the mean and median angular errors, where the lower values indicates better accuracy. We also measure the percentage of pixels with angular error below thresholds of $11.25^{\circ}$, $22.5^{\circ}$, and $30^{\circ}$. Quantitatively, SN and GeoWizard (GW), which are specifically designed for normal estimation tasks, outperform RGBX and DiffusionRender(DR), which are originally developed for image/video relighting. Visually RGBX and DR produce the surface normals with finer high-frequency details than SN and GW. However, none of these methods recover accurate surface normals for real-world objects with complex geometric structures in the evaluation dataset, highlighting the importance of OLATverse for advancing normal estimation research.
@@
Mitsuba+Mshape [nimier2019mitsuba] 35.906 0.0260 0.976
GS3 [bi2024gs3] 38.538 0.0263 0.982
RNG [fan2025rng] 32.065 0.051 0.962
BiGS [liu2024bigs] 32.98 0.0451 0.940
@@
SN [ye2024stablenormal] 31.85 30.25 8.93 34.00 55.40
RGBX [zeng2024rgb] 51.95 49.70 6.40 22.80 35.85
DR [liang2025diffusion] 34.88 33.28 8.13 31.00 50.15
GW [fu2024geowizard] 34.42 32.03 10.98 34.10 50.05
@@ 5 Limitation
Although we apply a linear polarized filter to cancel out the specular reflection during normal extraction, artifacts remain noticeable for objects with glossy materials or low-reflectance texture, consistent with the observation in RNHA [zhou2023relightable]. These artifacts arise from the weak signal-to-noise ratio of low-reflectance objects and the ambiguity between view-dependent reflection and surface normals. While the extracted surface normals and diffuse albedo are not exact ground truth, they can still provide valuable supervision signals for multi-modal training tasks. In addition, the ground truth meshes are not incorporated in our dataset due to hardware limitations. In the future, it would be interesting to integrate an advanced scanning system to jointly capture both appearance and geometry of real objects.